OpenStudy (anonymous):

sin2x=sinx, 0<=x<=2pi

hero (hero):

Hint: Subtract sin x from both sides to get \(\sin^2x - \sin x = 0\) Then factor out sin x common to both terms to get \(\sin x (\sin x - 1) =0\) Now use zero product property again.

OpenStudy (anonymous):

i think its sin2x not \[\sin ^{2}x\]

hero (hero):

Okay, so if it's \(\sin(2x)\), then that is equal to \(2\sin(x)\cos(x)\)

hero (hero):

Which means \(\sin(2x) = \sin x\) Subtract sin x from both sides to get \(\sin(2x) - \sin x = 0\) Replace sin(2x) with 2sin x cos x to get \(2\sin x \cos x - \sin x = 0\) Now factor out sin x common to both terms: \(\sin x(2\cos x-1) = 0\) Now use zero product property